# End-to-End Audio-Visual Learning for Cochlear Implant Sound Coding in Noisy Environments

### v1

@@ IV Results
ACE - 0.4870 0.2067 0.3262
ECS - 0.4870 0.2073 0.3258
ASE-ECS Pre-trained 0.5943 0.3557 0.4914
AVSE-ECS Pre-trained 0.6141 0.3672 0.5067
AVSE-ECS Joint training 0.6305 0.3899 0.5211
IV-A ECS with clean and noisy speech
The objective speech intelligibility of the ECS model is evaluated with noisy and clean speech, as shown in Table I. In the absence of noise, ECS achieves an STOI score of 0.7604. Under noisy conditions, the score drops to 0.4870. Similar reductions in ESTOI and NCM scores are observed as speech transitions from clean to noisy conditions. The result shows that the performance of the ECS model declines in the noisy condition, since ECS simulates the ACE strategy without an SE approach.
IV-B AVSE-ECS with different $\beta$ values
Table II presents the objective evaluation results of AVSE-ECS with different joint training setups. Under noisy conditions, AVSE-ECS achieves an STOI score of 0.6295 ($\beta$ = 0.25), 0.6305 ($\beta$ = 0.5), and 0.6196 ($\beta$ = 1).
Table II shows similar trends in STOI and ESTOI scores with the three $\beta$ values. These results suggest that emphasizing $\beta$, the weight on the electrodogram loss, by setting it to 1 does not necessarily improve speech intelligibility. On the other hand, a smaller $\beta$ value of 0.25 limits the refinement of the electrodogram. Among the tested values, $\beta$ = 0.5 results in the highest STOI and ESTOI scores among the three weights. Although $\beta=1$ achieves a slightly higher NCM score (0.5217) compared to $\beta=0.5$ (0.5211), the difference is marginal and not considered significant. Therefore, the $\beta$ value of 0.5 is selected for the subsequent experiment.
IV-C Comparison of ACE, ECS, ASE-ECS, and AVSE-ECS
@@
Table III compares different CI system implementations under noisy conditions. The ECS approach achieves a STOI score of 0.4870, which is comparable to that the ACE strategy. Using the pre-trained setup, the audio-only ASE-ECS model achieves a STOI score of 0.5943, surpassing the ECS strategy without the SE module. With the addition of visual cues, the AVSE-ECS model increases the STOI score to 0.6141 in the pre-trained setting and to 0.6305 with joint training. Similar trends can be observed in the ESTOI and NCM scores.
@@
These results demonstrate that the ECS model effectively emulates the performance of ACE, validating the feasibility of neural network–based CI systems. Furthermore, incorporating the ASE and AVSE models improves objective speech intelligibility compared to the ECS model. The use of visual cues in the AVSE-ECS model further enhances the SE capability compared to the audio-only ASE-ECS model. However, cleaner speech input does not necessarily result in CI electrode stimulation patterns with higher speech intelligibility. To address this, the proposed joint training strategy incorporates the clean electrodogram as an auxiliary training target. This additional training objective allows the AVSE model to produce enhanced speech that enables the CI system to generate electrode stimulation patterns with higher speech intelligibility. Therefore, with the joint training method, the AVSE-ECS model achieves the highest scores in this study.
IV-D Discussion
@@
A few aspects can be discussed for this study. While comprehensive objective evaluations have been conducted, additional subjective listening tests are planned to assess the effectiveness of the proposed AVSE-ECS method from a human perceptual perspective. These subjective experiments are to be conducted with both normal-hearing individuals using CI simulations and actual CI users. Physiological evaluations, including responses to auditory stimuli processed by the proposed approaches, may provide objective observations in human subjects.
@@
In this study, the objective evaluation relies on a single TMSV dataset in one language, Mandarin Chinese. As highlighted in prior studies [32, 33, 34], regression-based tasks such as speech enhancement and speech separation are generally less sensitive to language differences. As a result, models trained on English speech data have shown effective generalization to other languages. Moreover, to ensure the cross-language, cross-speaker, cross-dataset, and cross-validation generalizability of the proposed approach, further investigations using diverse datasets are planned. Since audiovisual datasets are less accessible than audio-only datasets, further dataset collection and development are needed in terms of size, diversity, and real-world coverage. For each dataset, multiple runs with different random seeds will be repeated to ensure the consistency and robustness of the proposed system.
@@
Implementing AVSE processing in CI hardware poses practical challenges, especially in achieving real-time performance. The proposed AVSE system with end-to-end CI processing can be executed on a PC, making it suitable for applications with relaxed timing constraints, such as movie streaming, online courses, and playback scenarios. Real-time deployment on CI sound processors is challenging due to latency, but the system is expected to integrate with external edge devices like smartphones, tablets, or smart glasses as intermediary platforms for practical use. These devices can offload computation from the implant-side processors, enabling more advanced audiovisual processing in everyday use. Looking ahead, we plan to further reduce the model’s computational footprint through techniques such as model pruning and knowledge distillation. In addition, further investigations into the NCSN++ model, the proposed AVSE backbone, will be conducted to assess the feasibility of using lightweight alternatives for hardware implementation and their seamless integration into the hearing assistance ecosystem. In the future, advances in device processing power are expected to reduce signal processing latency and make the AVSE module a feasible application.
@@
It is worth noting to explore the end-to-end architecture. A recent study demonstrates an analogous architecture to an end-to-end neural amplifier designed for hearing aids. [35]. Future research will aim to improve the overall CI sound processing system by exploring more advanced AVSE pre-processing methods and end-to-end frameworks.
@@ V Conclusion
This paper introduces a novel end-to-end CI system, AVSE-ECS, which integrates an audio-visual speech enhancement (AVSE) model with a deep-learning-based sound coding strategy, ElectrodeNet-CS (ECS). This system exhibits strong noise suppression capabilities and effectively mitigates severe performance degradation of the CI in noisy environments. A joint training approach used to train the entire end-to-end system enhances electrode stimulation patterns, making them more distinct and recognizable. A series of experiments is conducted to verify the effectiveness of the proposed method. Future work includes subjective listening evaluations, cross-dataset generalization studies, and exploration of alternative AVSE preprocessing frameworks. The methods and findings of this study may provide insights into end-to-end approaches for CI sound coding strategies based on multimodal processing, such as audio-visual integration.

### v2

@@ IV.2 AVSE-ECS with different $\beta$ values
ACE - 0.4870 0.2067 0.3262 -
ECS - 0.4870 0.2073 0.3258 0.0761
ASE-ECS Pre-trained 0.5943 0.3557 0.4914 3.8075
AVSE-ECS Pre-trained 0.6141 0.3672 0.5067 3.7291
AVSE-ECS Joint training 0.6305 0.3899 0.5211 7.4666
IV.3 Comparison of ACE, ECS, ASE-ECS, and AVSE-ECS
@@
Table 3 compares various CI system implementations under noisy conditions, with objective evaluation scores averaged across speakers, noise types and SNR levels. The ECS model achieves an average STOI score of 0.4870, comparable to the ACE strategy. With the pre-trained setup, the audio-only ASE-ECS model achieves a STOI score of 0.5943, surpassing the ECS strategy without the SE module. Incorporating visual cues in the AVSE-ECS model further increases STOI scores to 0.6141 (pre-trained) and 0.6305 (joint training). Similar trends appear in ESTOI and NCM scores.
The average SER scores for all models are higher than that of the ACE strategy. The SER difference for the ECS model compared to ACE is 0.0761 dB, while the scores for ASE-ECS (3.8075 dB) and pre-trained AVSE-ECS (3.7291 dB) are close to each other. Moreover, the joint-training model achieves the highest SER score, exceeding ACE by 7.4666 dB.
@@
Therefore, integrating ASE and AVSE models improves objective speech intelligibility over the ECS model, while the AVSE-ECS model with visual cues outperforms the audio-only ASE-ECS model in STOI, ESTOI, and NCM scores. In joint training that incorporates the clean electrodogram as an auxiliary target, the AVSE model can directly generate enhanced electrode stimulation patterns and achieve the highest scores in this study.
IV.4 Discussion
@@
This study can be further extended. Beyond the comprehensive objective evaluations, subjective listening tests are planned for the proposed AVSE-ECS method with both normal-hearing participants using CI simulations and actual CI users. Additionally, physiological evaluations, such as responses to auditory stimuli processed by the proposed approaches, may provide further objective evidence in human subjects.
@@
In this study, the objective evaluation relies on a single TMSV dataset in one language, Mandarin Chinese. As highlighted in prior studies [TTS_languate_zhang2019, TTS_languate_sankar2024, Mamba_chao2025], regression-based tasks such as speech enhancement and speech separation are generally less sensitive to language differences. As a result, models trained on English speech data have shown effective generalization to other languages. Moreover, to ensure the cross-language, cross-speaker, cross-dataset, and cross-validation generalizability of the proposed approach, further studies using diverse datasets are planned. Analyses with a wider range of noise types are to be conducted, as the present results indicate no clear differences among the five noise types examined. Since audiovisual datasets are less accessible than audio-only ones, further dataset collection and development are needed in terms of size, diversity, and real-world coverage. For each dataset, multiple runs with different random seeds will be repeated to ensure the consistency and robustness of the proposed system.
@@
Issues in implementing AVSE in CI hardware need to be tackled. The proposed AVSE system with end-to-end CI processing can be executed on a personal computer, making it suitable for applications with relaxed timing constraints, such as movie streaming, online courses, and playback scenarios. Latency is challenging for real-time deployment on CI sound processors, but integration with external edge devices such as smartphones, tablets, or smart glasses can offload computation from implant-side processors, enabling more advanced audiovisual processing in everyday use. Looking ahead, we plan to reduce the model’s computational load through techniques such as pruning and knowledge distillation. Further investigations into the NCSN++ model, the proposed AVSE backbone, will assess lightweight alternatives for hardware implementation and seamless integration into the hearing assistance ecosystem. In the future, advances in device processing power are expected to reduce signal processing latency and make the AVSE module a feasible application.
@@
It is worth noting to consider a number of related topics. A recent study demonstrates an analogous architecture to an end-to-end neural amplifier designed for hearing aids [NeuroAMP]. Therefore, more advanced AVSE pre-processing methods and end-to-end frameworks will be explored to enhance the overall CI sound processing system. Furthermore, the role of multisensory cues in CI perception will be investigated to provide more natural sound and enhance communication and well-being for CI users. Research on deep learning and CI sound processing is an ongoing journey.
@@ V Conclusion
This paper introduces a novel end-to-end CI system, AVSE-ECS, which integrates an audio-visual speech enhancement (AVSE) model with a deep-learning-based sound coding strategy, ElectrodeNet-CS (ECS). This system exhibits strong noise suppression capabilities and effectively mitigates severe performance degradation of the CI in noisy environments. Joint training of the end-to-end system enhances electrode stimulation patterns for greater clarity and recognizability, as verified through experiments. Future work includes subjective listening evaluations, cross-dataset generalization studies, and exploration of alternative AVSE preprocessing frameworks. The methods and findings of this study may provide insights into end-to-end CI coding strategies, multimodal processing, and related fields.
VI Author Declarations
VI.1 Conflict of Interest
The authors have no conflicts to disclose.
VI.2 Ethics Approval
This study was based solely on CI simulations and did not involve human participants, patient data, or animal subjects.
VII Data Availability
The data supporting the results of this research are available from the corresponding author upon reasonable request.